OpenStudy (anonymous):

7/8+3/4 X 1/4 how do I do this?

OpenStudy (anonymous):

OK Would that be \[\frac{ 7 }{ 8}+\frac{ 3 }{4 } *\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

OK First change 3/4 and 1/4 to have a base of 8

OpenStudy (anonymous):

Note since 4 is half of 8 the eaisest is to double them

OpenStudy (anonymous):

I.E 1/4 becomes 2/8

OpenStudy (anonymous):

Do the same for 3/4

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so is it 6/8?

OpenStudy (anonymous):

OK! So that gives you \[\frac{ 7 }{ 8 }+\frac{ 6 }{8 }*\frac{ 2 }{8 }\]

OpenStudy (anonymous):

15/8?

OpenStudy (anonymous):

oh wait nevermind lol

OpenStudy (texaschic101):

you do not have to change the 1/4

OpenStudy (anonymous):

I'm a little rusty on fractions sorry about that. Haven't done them manually for a few years

OpenStudy (anonymous):

so wait. now im confused

OpenStudy (anonymous):

OK 6/8*2/8?

OpenStudy (anonymous):

Times the top and bottom across 6*2=12 8*8=?

OpenStudy (anonymous):

64

OpenStudy (texaschic101):

do what jo said....add 7/8 + 6/8 and then multiply that answer by 1/4

OpenStudy (anonymous):

OK so now you have \[\frac{ 12 }{64 }\rightarrow \frac{ 3 }{ ? }\]

OpenStudy (anonymous):

sorry 3/16

OpenStudy (texaschic101):

or you can multiply it by 2/8....it doesn't matter, they are equivalent

OpenStudy (anonymous):

okay, so the answer would be 3/16?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

3/16 + 6/8

OpenStudy (anonymous):

im so confused. honestly. I don't get this. do you have to have common denominators to add that?

OpenStudy (texaschic101):

7/8 + 6/8 * 1/4 13/8 * 1/4 = 13/32

OpenStudy (debbieg):

You need an LCD for the addition step. NOT for the multiplication step.

OpenStudy (anonymous):

Yes so change 6/8 to 12/16

OpenStudy (anonymous):

so 15/16

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

I think i'm getting my algebra mixed up with your fractions

OpenStudy (texaschic101):

no.......7/8 + 6/8 * 1/4 13/8 * 1/4 = 13/32

OpenStudy (anonymous):

so its 13/32?

OpenStudy (anonymous):

well i'll just come straight out and tell you my calculator (given your origanal equation) is 17/16

OpenStudy (debbieg):

\[\frac{ 7 }{ 8 }+\frac{ 3 }{ 4 }\times\frac{ 1 }{ 4 }=\frac{ 7 }{ 8 }+\frac{ 3 }{ 16 }\]

OpenStudy (texaschic101):

kinda depends.... is your question : (7/8 + 3/4) * 1/4 or 7/8 + (3/4 * 1/4)

OpenStudy (anonymous):

I guess the first one. theres no parentheses in the equation tho

OpenStudy (anonymous):

What DebbieG said

OpenStudy (anonymous):

My mistake @jbrocksmysocks96

OpenStudy (debbieg):

Then you need LCD of 8 and 16, build up the first equation: \[=\frac{ 14 }{ 16 }+\frac{ 3 }{ 16 }=\frac{ 17 }{ 16 }\]

OpenStudy (anonymous):

Thanks @DebbieG !

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

thank you !!

OpenStudy (debbieg):

no problem. :)

OpenStudy (anonymous):

Been Awhile since I've done fractions!

OpenStudy (anonymous):

yeah, that's why im confused I have a couple more but I don't wanna bother you

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (anonymous):

As in one that does fractions?

OpenStudy (texaschic101):

I guess according to order of operations, you would multiply first

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

You can get one for around 20-30 dollars

OpenStudy (anonymous):

well worth it

OpenStudy (anonymous):

okay. well I need to figure out these questions now. I don't know what to do. I really don't understand fractions period

OpenStudy (texaschic101):

fractions are not that hard......when adding/subtracting fractions you need a common denominator. When multiplying, you just multiply straight across. When dividing, you " flip " the number you are dividing by, then multiply

OpenStudy (anonymous):

Just remember this To add make sure the denominator (base number) are the same and add the top number To multiply just multiply both the top and bottom number across.

OpenStudy (anonymous):

And remember PEMDAS

OpenStudy (texaschic101):

good point :)

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

PLEASE (PARENTHESIS) EXCUSE (EXPONENTS) MY (MULTIPLY) DEAR (DIVIDE) AUNT (ADD) SALLY (SUBTRACT and multiplying and dividing and adding and subtracting are interchangeable if dividing comes becofe multiplying (or vice versa) do which ever comes first

OpenStudy (texaschic101):

you need anymore help ?

OpenStudy (anonymous):

So Parenthesis first then exponents then multiplying &dividing or vice versa then adding & subtracting or vice versa

OpenStudy (anonymous):

this one equation is confusing me. (3/5-4/9) divided by 2/3

OpenStudy (debbieg):

And when possible, always REDUCE first, THEN multiply. It will save you work reducing after the fact!

OpenStudy (texaschic101):

(3/5 - 4/9) / (2/3) do the subtraction first...find the common denominator of 5 and 9

OpenStudy (anonymous):

OK \[(\frac{ 3 }{ 5 }-\frac{4}{9})/\frac{ 2 }{3 }\]

OpenStudy (anonymous):

PEDMAS do what's in the parenthesis first common demenonater of 5 &9?

OpenStudy (anonymous):

45? I don't know :(

OpenStudy (anonymous):

RIGHT!

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

So \frac{3}{5} \frac{4} {9} to a base of 45

OpenStudy (anonymous):

Sorry 3/5 & 4/9 change that to where the denominator is 45

OpenStudy (anonymous):

so I multiply everything in 3/5 by 9? to get 27/45?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Do the same for 4/9

OpenStudy (anonymous):

20/45

OpenStudy (anonymous):

and then id be left with 7/45 right?

OpenStudy (anonymous):

Great! Now subtract \[\frac{ 27 }{ 45 }-\frac{ 20 }{ 45 }\]

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

so now divide 7/45 by 2/3?

OpenStudy (anonymous):

Do you know how to?

OpenStudy (anonymous):

don't I have to do the reciprocal?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

so 7/45 times 3/2?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

21/90?

OpenStudy (anonymous):

now simplify and your done!

OpenStudy (anonymous):

7/30?

OpenStudy (anonymous):

Correcto!

OpenStudy (anonymous):

okay. I have one last one and that's it. if i explain each step can you go over it with me?

OpenStudy (anonymous):

I can try :)

OpenStudy (anonymous):

3/4-1/2 divided by 6/7

OpenStudy (anonymous):

so first find a common denonminator of 4 and 2?

OpenStudy (anonymous):

\[\frac{ 3 }{4}-\frac{ 1 }{ 2}/\frac{ 6 }{7 }\]

OpenStudy (anonymous):

yes that's it

OpenStudy (anonymous):

Not yet! Remember PEDMAS?

OpenStudy (anonymous):

oh divide first

OpenStudy (anonymous):

Thre arn't any parnthesis around 3/4 and 1/2 is there?